# How weak are weak factors? Uniform inference for signal strength in signal plus noise models

### v1

@@ 9.4. Canonical Correlation Analysis
The proof of Theorem 4.6 for CCA of Section 2.4 follows the same outline as in the previous three sections. However, an analogue of Propositions 9.1, 9.6, 9.10 becomes even more complicated, which introduces new challenges in the proofs. Note that the CCA setting is symmetric under $N\leftrightarrow M$ and there is no loss of generality to assume $N\leq M$, as we do throughout this section. The reader may consult Bykhovskaya and Gorin (2024) for general information on CCA.
Suppose that we are given $N\times S$ matrix $\mathbf{U}$ and $M\times S$ matrix $\mathbf{V}$, in which the first rows are denoted $\mathbf{u}^{*}$ and $\mathbf{v}^{*}$, respectively. The remaining rows form matrices $\widetilde{\mathbf{U}}$ and $\widetilde{\mathbf{V}}$ of sizes $(N-1)\times S$ and $(M-1)\times S$, respectively. We would like to connect sample canonical correlations between $\mathbf{U}$ and $\mathbf{V}$ to $\mathbf{u}^{*}$, $\mathbf{v}^{*}$, and the sample canonical correlations and variables between $\widetilde{\mathbf{U}}$ and $\widetilde{\mathbf{V}}$. The latter are denoted $c_{1},\dots,c_{N-1}$, $\mathbf{u}_{1},\dots,\mathbf{u}_{N-1}$, $\mathbf{v}_{1},\dots,\mathbf{v}_{M-1}$, where $1\geq c_{1}\geq\dots\geq c_{N-1}\geq 0$, $\mathbf{u}_{i}$ are $S$–dimensional vectors forming an orthonormal basis in the space spanned by rows of $\widetilde{\mathbf{U}}$, $\mathbf{v}_{j}$ are $S$–dimensional vectors forming an orthonormal basis in the space spanned by the rows of $\widetilde{\mathbf{V}}$, and $\langle\mathbf{u}_{i},\mathbf{v}_{j}\rangle=c_{i}\delta_{i=j}$, $11\leq i\leq N-$ 1, $11\leq j\leq M-$ 1. For the convenience of notation, we also introduce numbers $c_{N}=c_{N+1}=\dots=c_{M-1}=0$.
@@
(9.54) $\begin{split}&\Bigg{[}\langle\mathbf{u}^{*},\mathbf{v}^{*}\rangle+\sum_{j=1}^{M-1}\frac{\langle\mathbf{u}^{*},\mathbf{v}_{j}\rangle(c_{j}\langle\mathbf{v}^{*},\mathbf{u}_{j}\rangle-a\langle\mathbf{v}^{*},\mathbf{v}_{j}\rangle)}{a-c_{j}^{2}}\\ &\qquad\qquad-a\sum_{i=1}^{N-1}\frac{\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle(\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle-c_{i}\langle\mathbf{v}^{*},\mathbf{v}_{i}\rangle)}{a-c_{i}^{2}}\Bigg{]}^{2}\\ &\qquad=a\left[-\langle\mathbf{u}^{*},\mathbf{u}^{*}\rangle+\sum_{j=1}^{M-1}\frac{\langle\mathbf{u}^{*},\mathbf{v}_{j}\rangle^{2}-2c_{j}\langle\mathbf{u}^{*},\mathbf{v}_{j}\rangle\langle\mathbf{u}^{*},\mathbf{u}_{j}\rangle}{a-c_{j}^{2}}+a\sum_{i=1}^{N-1}\frac{\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle^{2}}{a-c_{i}^{2}}\right]\\ &\qquad\qquad\times\left[-\langle\mathbf{v}^{*},\mathbf{v}^{*}\rangle+\sum_{i=1}^{N-1}\frac{\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle^{2}-2c_{i}\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle\langle\mathbf{v}^{*},\mathbf{v}_{i}\rangle}{a-c_{i}^{2}}+a\sum_{j=1}^{M-1}\frac{\langle\mathbf{v}^{*},\mathbf{v}_{j}\rangle^{2}}{a-c_{j}^{2}}\right],\end{split}$
@@
We omit the proof, see Bykhovskaya and Gorin (2025, Appendix A), with $N=K$. The next equation replaces (9.8), (9.19), (9.38) in the context of CCA.
@@
where $(\xi_{i},\eta_{i})$ are i.i.d. Gaussian mean 00 two-dimensional vectors with covariance matrix
@@
We fix the index $1\leq k\leq r$ and connect Proposition 9.15 to the setting of Section 2.4. For that we choose two deterministic invertible matrices $\widetilde{A}$ of size $N\times N$ and $\widetilde{B}$ of size $M\times M$, so that:
@@
The existence of such $\widetilde{A}$ and $\widetilde{B}$ readily follows from the existence of the decomposition (2.14), which is the basic statement in CCA – the first coordinates of $\widetilde{A}\mathbf{u}$ and $\widetilde{B}\mathbf{v}$ are canonical variables, corresponding to the canonical correlation $\theta_{k}$.
Multiplying the matrices $\mathbf{U}$ and $\mathbf{V}$ by $\widetilde{A}$ and $\widetilde{B}$, respectively, does not change the squared sample canonical correlations, and brings them to the form of Proposition 9.15. It remains to explain that (9.54) is the same as (9.55). The correlation structure between the last $N-1$ rows of $\widetilde{A}\mathbf{U}$ and the last $M-1$ rows of $\widetilde{B}\mathbf{V}$ is as in (2.14) but with $\theta_{k}$ removed from $\{\theta_{1},\theta_{2},\dots,\theta_{r}\}$. Hence $c_{i}^{2}$ in (9.54) match their description in Corollary 9.16. The first rows of $\widetilde{A}\mathbf{U}$ and $\widetilde{B}\mathbf{V}$ are $S$–dimensional vectors with i.i.d. components, and with correlation structure as in (9.56). We should further understand the joint distribution of the four arrays of random scalar products appearing in (9.54).
@@
If all $\mathbf{u}_{i}$,$\mathbf{v}_{j}$ were orthonormal, then computing scalar products would be easy: scalar products of i.i.d. mean 00 Gaussian vectors with orthonormal basis are again i.i.d. Gaussian vectors. However, $\langle\mathbf{u}_{i},\mathbf{v}_{i}\rangle\neq 0$. In order to fix this, we introduce new vectors $\mathbf{v}^{\prime}_{j}$, $11\leq j\leq M-$ 1, which are:
@@
Since $\{\mathbf{u}_{i}\}$ are orthonormal, $\{\mathbf{v}_{j}\}$ are orthonormal, and $\langle\mathbf{u}_{i},\mathbf{v}_{j}\rangle=c_{i}\delta_{i=j}$, we conclude that the $N+M-2$ vectors $\mathbf{u}_{1},\dots,\mathbf{u}_{N-1}$, $\mathbf{v}^{\prime}_{1},\dots,\mathbf{v}^{\prime}_{M-1}$ are also orthonormal.
Since $\mathbf{u}^{*},\mathbf{v}^{*}$ are Gaussian and independent from all $\mathbf{u}_{j}$, $\mathbf{v}_{j}$, the scalar products $\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle,\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle,\langle\mathbf{u}^{*},\mathbf{v}^{\prime}_{j}\rangle,\langle\mathbf{v}^{*},\mathbf{v}^{\prime}_{j}\rangle$ are also Gaussian and independent from all $\mathbf{u}_{j}$ and $\mathbf{v}_{j}$. Among these scalar products, most are pairwise independent, with the only non-zero covariances being:
@@
Let us choose an orthonormal basis $\mathbf{w}_{1},\dots,\mathbf{w}_{S}$ of the $S$–dimensional space, such that the first $N+M-2$ vectors are $\mathbf{u}_{1},\dots,\mathbf{u}_{N-1}$, $\mathbf{v}^{\prime}_{1},\dots,\mathbf{v}^{\prime}_{M-1}$ and the rest are arbitrary (the choice is independent from $\mathbf{u}^{*}$ and $\mathbf{v}^{*}$). Define
$\xi_{i}=\langle\mathbf{u}^{*},\mathbf{w}_{i}\rangle,\qquad\eta_{i}=\langle\mathbf{v}^{*},\mathbf{w}_{i}\rangle.$
Then $(\xi_{i},\eta_{i})$, $1\leq i\leq S$, are i.i.d. vectors with correlation structure (9.56). We express various scalar products in (9.54) trough $\xi_{i}$ and $\eta_{i}$:
$\langle\mathbf{u}^{*},\mathbf{v}^{*}\rangle=\sum_{i=1}^{S}\xi_{i}\eta_{i},\qquad\langle\mathbf{u}^{*},\mathbf{u}^{*}\rangle=\sum_{i=1}^{S}\xi_{i}^{2},\qquad\langle\mathbf{v}^{*},\mathbf{v}^{*}\rangle=\sum_{i=1}^{S}\eta_{i}^{2}.$
$\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle=\xi_{i},\,\,\,\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle=\eta_{i},$
$\langle\mathbf{u}^{*},\mathbf{v}_{j}\rangle=\sqrt{1-c_{j}^{2}}\,\xi_{j+N-1}+c_{j}\xi_{j},\,\,\,\langle\mathbf{v}^{*},\mathbf{v}_{j}\rangle=\sqrt{1-c_{j}^{2}}\,\eta_{j+N-1}+c_{j}\eta_{j}.$
Plugging these expressions into (9.54), we get (9.55). ∎
We record a version of Corollaries 9.3, 9.8, 9.12, which is slightly more complicated, leading to an additional step in the proofs, on which we comment in Remark 9.20.
@@
Let $N$ roots of (9.54) be denoted $a_{1},\dots,a_{N}$. Then all $a_{i}$ are real numbers between 00 and $1$. Moreover, if we arrange $a_{i}$ in the decreasing order, then there exists another sequence of ${N-1}$ real numbers $y_{1}\geq y_{2}\geq\dots\geq y_{{N-1}}$, such that two interlacing conditions hold:
@@
In the notation of the proof of Corollary 9.16, the numbers $y_{1},\dots,y_{N-1}$ are squared sample canonical correlations between $\widetilde{\mathbf{U}}=$(last $N-1$ rows of $\widetilde{A}\mathbf{U}$) and $\mathbf{V}$
@@
This is a version of Lemma A.6 in Bykhovskaya and Gorin (2025). $a_{1},\dots,a_{N}$ are eigenvalues of the product of projections on $\mathbf{U}$ and on $\mathbf{V}$ in $S$–dimensional space, i.e., $P_{\mathbf{U}}P_{\mathbf{V}}$ (or $P_{\mathbf{V}}P_{\mathbf{U}}P_{\mathbf{V}}$ or $P_{\mathbf{U}}P_{\mathbf{V}}P_{\mathbf{U}}$), while $y_{1},\dots,y_{N-1}$ are eigenvalues of the product of projections on smaller $\widetilde{\mathbf{U}}$ and $\mathbf{V}$. Then (9.58) are two instances interlacing inequalities between eigenvalues of a matrix and its principal submatrix, as in Bhatia (1997, Corollry III.1.5). ∎
Next, we prove an analogue of Propositions 9.5, 9.9, 9.14.
@@
Consider the CCA model, as in Section 2.4, with $\theta_{1}>\theta_{2}>\dots>\theta_{r}\geq 0$ split into two groups: $\theta_{1},\dots,\theta_{q-1}>\theta^{c}=\frac{1}{\sqrt{(\tau_{M}-1)(\tau_{N}-1)}}$ and $\theta_{q},\dots,\theta_{r}<\theta^{c}=\frac{1}{\sqrt{(\tau_{M}-1)(\tau_{N}-1)}}$. Then, with $\frac{S}{N}=\tau_{N}+O\left(\tfrac{1}{N}\right),\quad\frac{S}{M}=\tau_{M}+O\left(\tfrac{1}{N}\right)$, in the sense of convergence in joint distribution and using (2.16):
@@
where $\mathcal{N}(0,V(\theta_{i}))$, $10\leq i\leq q-$ 1, are independent and $\{\mathfrak{a}_{j}\}_{j\geq 1}$ are points of the Airy1 point process independent from (9.59). In addition, (8.46) hold for the sample squared canonical correlations between $\mathbf{U}$ and $\mathbf{V}$ with $h(x)=1$, $h(x)=\sqrt{x}$, $h(x)=\sqrt{1-x}$, and $h(x)=\sqrt{x(1-x)}$.
@@
Statements close to (9.59), (9.60) are known from Bao et al. (2019), Yang (2022b). Alternatively, the proof can be obtained by the same argument as in Proposition 9.5, by induction on $r$ with the base case $r=0$ given in Johnstone (2008); Han et al. (2018) and the step based on Corollary 9.16. We only highlight the key computation, which is an analogue of (9.11), (9.12).
We recall (9.55) and analyze its asymptotic behavior for $a>\lambda_{+}$ and assuming that $\{c_{i}^{2}\}$ are squared sample canonical correlations between independent $\mathbf{U}$ and $\mathbf{V}$ of sizes $(N-1)\times S$ and $(M-1)\times S$, respectively, and with i.i.d. $\mathcal{N}(0,1)$ random variables as their elements. In this situation the joint distribution of $\{c_{i}^{2}\}$ is the JOE ensemble (see Section 8.4), and their empirical measure converges to the Wachter law (9.61) as $N\to\infty$. According to Remark 9.17, we use (9.56) with $C_{uu}=C_{vv}=1$, $C_{uv}=\sqrt{\theta_{k}}$; we also recall $c_{j}=0$ for $j\geq N$. We divide (9.55) by $N^{2}$ and split its sums into their expectations (conditional on $\{c_{i}\}$) and the mean 00 parts. The left-hand side of (9.55) is transformed into:
@@
where the first line is of constant order and deterministic as $N\to\infty$; the second and third lines are of order $N^{-1/2}$ and become Gaussian as $N\to\infty$. The []delimited-[][$\cdot$][ ⋅ ] factor in the third line of (9.55) is similarly transformed into:
@@
Further, similarly to the proofs of Propositions 9.5, 9.9, and 9.14, applying CLT to the second-order terms in (9.64), (9.65), (9.66), we replace $O(N^{-1/2})$ in (9.67) with $N^{-1/2}$ times $\mathcal{N}(0,1)$ times an explicit function of $a$. Then solving (9.67) again, we arrive at the statement on the asymptotic Gaussianity of the solution $a$, which matches (2.16). ∎
@@
In the proof of Proposition 9.5 we located all the largest eigenvalues by analyzing the equation (9.1) for the variable $a$ near each $\lambda(\theta_{i})$ and near $\lambda_{+}$. In addition, we used the interlacements of Corollary 9.3 to make sure that we did not miss any other large eigenvalues. Similarly in the above proof, by analyzing (9.55), we locate $a$ solving the equation near each $\lambda(\theta_{i})$ and near $\lambda_{+}$. The argument for not missing other eigenvalues needs to be modified this time: if we use (9.58) naïvely, then, say, in the $r=1$ case it would allow two eigenvalues $\lambda_{1}\geq\lambda_{2}$ larger than $c_{1}^{2}$, rather than only one in Corollary 9.3.
The remedy is to still use (9.58), but compare with $y_{1},\dots,y_{N-1}$ instead of $c_{1}^{2},\dots,c_{N}^{2}$. Note that in the induction step of the argument of Proposition 9.19, the largest eigenvalues $y_{1},y_{2},\dots$, satisfy exactly the same induction assumption and therefore exactly the same asymptotics as $c_{1}^{2},c_{2}^{2},\dots$ (the only difference is that $N$ got decreased by $1$, but this does not change the asymptotic behavior). With this update, the same arguments, as before, go through. The same remark applies to the following proof.
@@
(9.69) $\lambda^{\prime\prime}(\theta^{c})$ $=$ $2\frac{(\tau_{N}-1)^{3/2}(\tau_{M}-1)^{3/2}}{\tau_{N}\tau_{M}}\leavevmode\nobreak\ ,$

### v2

@@ 9.4. Canonical Correlation Analysis
The proof of Theorem 4.6 for CCA of Section 2.4 follows the same outline as in the previous three sections. However, an analogue of Propositions 9.1, 9.6, 9.11 becomes even more complicated, which introduces new challenges in the proofs. The setting is symmetric under $N\leftrightarrow M$ and we assume $N\leq M$. The reader may consult Bykhovskaya and Gorin (2024) for general information on CCA.
Suppose that we are given $N\times S$ matrix $\mathbf{U}$ and $M\times S$ matrix $\mathbf{V}$, in which the first rows are denoted $\mathbf{u}^{*}$ and $\mathbf{v}^{*}$, respectively. The remaining rows form matrices $\widetilde{\mathbf{U}}$ and $\widetilde{\mathbf{V}}$ of sizes $(N-1)\times S$ and $(M-1)\times S$, respectively. We would like to connect sample canonical correlations between $\mathbf{U}$ and $\mathbf{V}$ to $\mathbf{u}^{*}$, $\mathbf{v}^{*}$, and the sample canonical correlations and variables between $\widetilde{\mathbf{U}}$ and $\widetilde{\mathbf{V}}$. The latter are denoted $c_{1},\dots,c_{N-1}$, $\mathbf{u}_{1},\dots,\mathbf{u}_{N-1}$, $\mathbf{v}_{1},\dots,\mathbf{v}_{M-1}$, where $1\geq c_{1}\geq\dots\geq c_{N-1}\geq 0$, $\mathbf{u}_{i}$ are SS–dimensional vectors forming an orthonormal basis in the space spanned by rows of $\widetilde{\mathbf{U}}$, $\mathbf{v}_{j}$ are SS–dimensional vectors forming an orthonormal basis in the space spanned by the rows of $\widetilde{\mathbf{V}}$, and $\langle\mathbf{u}_{i},\mathbf{v}_{j}\rangle=c_{i}\delta_{i=j}$, $1\leq i\leq N-1$, $1\leq j\leq M-1$. For the convenience of notation, we also introduce numbers $c_{N}=c_{N+1}=\dots=c_{M-1}=0$.
@@
(9.54) $\begin{split}&\Bigg[\langle\mathbf{u}^{*},\mathbf{v}^{*}\rangle+\sum_{j=1}^{M-1}\frac{\langle\mathbf{u}^{*},\mathbf{v}_{j}\rangle(c_{j}\langle\mathbf{v}^{*},\mathbf{u}_{j}\rangle-a\langle\mathbf{v}^{*},\mathbf{v}_{j}\rangle)}{a-c_{j}^{2}}\,-a\sum_{i=1}^{N-1}\frac{\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle(\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle-c_{i}\langle\mathbf{v}^{*},\mathbf{v}_{i}\rangle)}{a-c_{i}^{2}}\Bigg]^{2}\\ &\qquad=a\left[-\langle\mathbf{u}^{*},\mathbf{u}^{*}\rangle+\sum_{j=1}^{M-1}\frac{\langle\mathbf{u}^{*},\mathbf{v}_{j}\rangle^{2}-2c_{j}\langle\mathbf{u}^{*},\mathbf{v}_{j}\rangle\langle\mathbf{u}^{*},\mathbf{u}_{j}\rangle}{a-c_{j}^{2}}+a\sum_{i=1}^{N-1}\frac{\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle^{2}}{a-c_{i}^{2}}\right]\\ &\qquad\qquad\times\left[-\langle\mathbf{v}^{*},\mathbf{v}^{*}\rangle+\sum_{i=1}^{N-1}\frac{\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle^{2}-2c_{i}\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle\langle\mathbf{v}^{*},\mathbf{v}_{i}\rangle}{a-c_{i}^{2}}+a\sum_{j=1}^{M-1}\frac{\langle\mathbf{v}^{*},\mathbf{v}_{j}\rangle^{2}}{a-c_{j}^{2}}\right],\end{split}$
@@
We omit the proof, see Bykhovskaya and Gorin (2025, Appendix A), with $N=K$.
@@
where $(\xi_{i},\eta_{i})$ are i.i.d. Gaussian mean 0 two-dimensional vectors with covariance matrix
@@
We fix $1\leq k\leq r$ and connect Proposition 9.16 to Section 2.4. We choose two deterministic invertible matrices $\widetilde{A}$ of size $N\times N$ and $\widetilde{B}$ of size $M\times M$, so that:
@@
The existence of such $\widetilde{A}$ and $\widetilde{B}$ is the basic statement in CCA, equivalent to (2.14) – the first coordinates of $\widetilde{A}\mathbf{u}$ and $\widetilde{B}\mathbf{v}$ are canonical variables for the canonical correlation $\theta_{k}$.
Multiplying the matrices $\mathbf{U}$ and $\mathbf{V}$ by $\widetilde{A}$ and $\widetilde{B}$, respectively, does not change the squared sample canonical correlations, and brings them to the form of Proposition 9.16. It remains to explain that (9.54) is the same as (9.55). The correlation structure between the last N1N-1 rows of $\widetilde{A}\mathbf{U}$ and the last M1M-1 rows of $\widetilde{B}\mathbf{V}$ is as in (2.14) but with $\theta_{k}$ removed from $\{\theta_{1},\theta_{2},\dots,\theta_{r}\}$. Hence $c_{i}^{2}$ in (9.54) match their description in Corollary 9.17. The first rows of $\widetilde{A}\mathbf{U}$ and $\widetilde{B}\mathbf{V}$ are SS–dimensional vectors with i.i.d. components, and with correlation structure as in (9.56). We should further understand the joint distribution of the four arrays of random scalar products appearing in (9.54).
@@
If all $\mathbf{u}_{i}$$\mathbf{v}_{j}$ were orthonormal, we could use that scalar products of i.i.d. mean 0 Gaussian vectors with orthonormal basis are again i.i.d. Gaussian vectors. However, $\langle\mathbf{u}_{i},\mathbf{v}_{i}\rangle\neq 0$. In order to fix this, we introduce new vectors $\mathbf{v}^{\prime}_{j}$, $1\leq j\leq M-1$, which are:
@@
Since $\{\mathbf{u}_{i}\}$ are orthonormal, $\{\mathbf{v}_{j}\}$ are orthonormal, and $\langle\mathbf{u}_{i},\mathbf{v}_{j}\rangle=c_{i}\delta_{i=j}$, we conclude that the N+M2N+M-2 vectors $\mathbf{u}_{1},\dots,\mathbf{u}_{N-1}$, $\mathbf{v}^{\prime}_{1},\dots,\mathbf{v}^{\prime}_{M-1}$ are also orthonormal.
Since $\mathbf{u}^{*},\mathbf{v}^{*}$ are Gaussian and independent from all $\mathbf{u}_{j}$, $\mathbf{v}_{j}$, the scalar products $\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle,\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle,\langle\mathbf{u}^{*},\mathbf{v}^{\prime}_{j}\rangle,\langle\mathbf{v}^{*},\mathbf{v}^{\prime}_{j}\rangle$ are also Gaussian and independent from all $\mathbf{u}_{j}$ and $\mathbf{v}_{j}$. Among these scalar products, he only non-zero covariances are:
@@
Let us choose an orthonormal basis $\mathbf{w}_{1},\dots,\mathbf{w}_{S}$ of the SS–dimensional space, such that the first N+M2N+M-2 vectors are $\mathbf{u}_{1},\dots,\mathbf{u}_{N-1}$, $\mathbf{v}^{\prime}_{1},\dots,\mathbf{v}^{\prime}_{M-1}$ and the rest are arbitrary (the choice is independent from $\mathbf{u}^{*}$ and $\mathbf{v}^{*}$). Define $\xi_{i}=\langle\mathbf{u}^{*},\mathbf{w}_{i}\rangle$, $\eta_{i}=\langle\mathbf{v}^{*},\mathbf{w}_{i}\rangle$. Then $(\xi_{i},\eta_{i})$, $1\leq i\leq S$, are i.i.d. vectors with correlation structure (9.56). We express various scalar products in (9.54) through $\xi_{i}$ and $\eta_{i}$: $\langle\mathbf{u}^{*},\mathbf{v}^{*}\rangle=\sum_{i=1}^{S}\xi_{i}\eta_{i},\quad\langle\mathbf{u}^{*},\mathbf{u}^{*}\rangle=\sum_{i=1}^{S}\xi_{i}^{2},\quad\langle\mathbf{v}^{*},\mathbf{v}^{*}\rangle=\sum_{i=1}^{S}\eta_{i}^{2};$ $\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle=\xi_{i},\,\,\,\langle\mathbf{v}^{*},\mathbf{u}_{i}\rangle=\eta_{i};$ $\langle\mathbf{u}^{*},\mathbf{v}_{j}\rangle=\sqrt{1-c_{j}^{2}}\,\xi_{j+N-1}+c_{j}\xi_{j},\,\,\,\langle\mathbf{v}^{*},\mathbf{v}_{j}\rangle=\sqrt{1-c_{j}^{2}}\,\eta_{j+N-1}+c_{j}\eta_{j}.$ Plugging these expressions into (9.54), we get (9.55). ∎
We record a version of Corollaries 9.3, 9.8, 9.13, which is slightly more complicated, leading to an additional step in the proofs, on which we comment in Remark 9.21.
@@
Let $N$ roots of (9.54) be denoted $a_{1},\dots,a_{N}$. Then all $a_{i}$ are real numbers between 0 and $1$. If we arrange $a_{i}$ in the decreasing order, then there exists another sequence of N1{N-1} real numbers $y_{1}\geq y_{2}\geq\dots\geq y_{{N-1}}$, such that two interlacing conditions hold:
@@
In the notation of the proof of Corollary 9.17, the numbers $y_{1},\dots,y_{N-1}$ are squared sample canonical correlations between $\widetilde{\mathbf{U}}$=(last N1N-1 rows of $\widetilde{A}\mathbf{U}$) and $\mathbf{V}$ .
@@
This is a version of Lemma A.6 in Bykhovskaya and Gorin (2025). $a_{1},\dots,a_{N}$ are eigenvalues of the product of projections on $\mathbf{U}$ and on $\mathbf{V}$ in SS–dimensional space, i.e., $P_{\mathbf{U}}P_{\mathbf{V}}$ (or $P_{\mathbf{V}}P_{\mathbf{U}}P_{\mathbf{V}}$ or $P_{\mathbf{U}}P_{\mathbf{V}}P_{\mathbf{U}}$), while $y_{1},\dots,y_{N-1}$ are eigenvalues of the product of projections on smaller $\widetilde{\mathbf{U}}$ and $\mathbf{V}$. Then (9.58) are two instances interlacing inequalities between eigenvalues of a matrix and its principal submatrix, as in Bhatia (1997, Corollry III.1.5). ∎
Next, we prove an analogue of Propositions 9.5, 9.9, 9.15.
@@
Consider the CCA model, with $\theta_{1}>\theta_{2}>\dots>\theta_{r}\geq 0$ split into $\theta_{1},\dots,\theta_{q-1}>\theta^{c}=\frac{1}{\sqrt{(\tau_{M}-1)(\tau_{N}-1)}}$ and $\theta_{q},\dots,\theta_{r}<\theta^{c}=\frac{1}{\sqrt{(\tau_{M}-1)(\tau_{N}-1)}}$. Then, with $\frac{S}{N}=\tau_{N}+O\left(\tfrac{1}{N}\right)$, $\frac{S}{M}=\tau_{M}+O\left(\tfrac{1}{N}\right)$, using (2.16), in joint distribution we have:
@@
where $\mathcal{N}(0,V(\theta_{i}))$ are independent over $i$ and with points of the Airy1 point process $\{\mathfrak{a}_{j}\}_{j\geq 1}$. In addition, (8.45) hold for the sample squared canonical correlations between $\mathbf{U}$ and $\mathbf{V}$ with $h(x)=1$, $h(x)=\sqrt{x}$, $h(x)=\sqrt{1-x}$, and $h(x)=\sqrt{x(1-x)}$.
@@
Statements close to (9.59), (9.60) are known from Bao et al. (2019), Yang (2022b). Alternatively, the proof can be obtained by the argument of Proposition 9.5, by induction on $r$ with case $r=0$ given in Johnstone (2008); Han et al. (2018) and the step based on Corollary 9.17. We only highlight the key computation, which is an analogue of (9.11), (9.12).
We recall (9.55) and analyze its asymptotic behavior for $a>\lambda_{+}$ and assuming that $\{c_{i}^{2}\}$ are squared sample canonical correlations between independent $\mathbf{U}$ and $\mathbf{V}$ of sizes $(N-1)\times S$ and $(M-1)\times S$, respectively, and with i.i.d. $\mathcal{N}(0,1)$ random variables as their elements. In this situation the joint distribution of $\{c_{i}^{2}\}$ is the JOE ensemble (see Section 8.4), and their empirical measure converges to the Wachter law (9.61) as $N\to\infty$. According to Remark 9.18, we use (9.56) with $C_{uu}=C_{vv}=1$, $C_{uv}=\sqrt{\theta_{k}}$; we also recall $c_{j}=0$ for $j\geq N$. We divide (9.55) by $N^{2}$ and split its sums into their expectations (conditional on $\{c_{i}\}$) and the mean 0 parts. The left-hand side of (9.55) is transformed into:
@@
where the first line is of constant order and deterministic as $N\to\infty$; the second and third lines are of order $N^{-1/2}$ and become Gaussian as $N\to\infty$. The $\cdot$] factor in the third line of (9.55) is similarly transformed into:
@@
Further, similarly to the proofs of Propositions 9.5, 9.9, and 9.15, applying CLT to the second-order terms in (9.64), (9.65), (9.66), we replace $O(N^{-1/2})$ in (9.67) with $N^{-1/2}$ times $\mathcal{N}(0,1)$ times an explicit function of $a$. Then solving (9.67) again, we arrive at the statement on the asymptotic Gaussianity of the solution $a$, which matches (2.16). ∎
@@
In Proposition 9.5 we located the largest eigenvalues by analyzing (9.1) for $a$ near each $\lambda(\theta_{i})$ and near $\lambda_{+}$. We used the interlacements of Corollary 9.3 to make sure that we did not miss any eigenvalues. Similarly in the above proof, by analyzing (9.55), we can locate solution $a$ near each $\lambda(\theta_{i})$ and near $\lambda_{+}$. The argument for not missing other eigenvalues needs to be modified: if we use (9.58) naïvely, then, say, in the $r=1$ case it would allow two eigenvalues $\lambda_{1}\geq\lambda_{2}$ larger than $c_{1}^{2}$, rather than only one in Corollary 9.3.
We still use (9.58), but compare with $y_{1},\dots,y_{N-1}$ instead of $c_{1}^{2},\dots,c_{N}^{2}$. In the induction step of the argument, the eigenvalues $y_{1},y_{2},\dots$, satisfy the same induction assumption and the same asymptotics as $c_{1}^{2},c_{2}^{2},\dots$ (the only difference is that $N$ got decreased by $1$). With this update, the same arguments, as before, go through. The same applies below.
@@
(9.69) $\lambda^{\prime\prime}(\theta^{c})$ $=$ $2\frac{(\tau_{N}-1)^{3/2}(\tau_{M}-1)^{3/2}}{\tau_{N}\tau_{M}}\penalty 10000\ ,$